### v2

@@ Abstract
Embodied artificial intelligence (EAI) integrates advanced AI models into physical entities for real-world interaction. The emergence of foundation models as the “brain” of EAI agents for high-level task planning has shown promising results. However, the deployment of these agents in physical environments presents significant safety challenges. For instance, a housekeeping robot lacking sufficient risk awareness might place a metal container in a microwave, potentially causing a fire. To address these critical safety concerns, comprehensive pre-deployment risk assessments are imperative. This study introduces EAIRiskBench , a novel framework for automated physical risk assessment in EAI scenarios. EAIRiskBench employs a multi-agent cooperative system that leverages various foundation models to generate safety guidelines, create risk-prone scenarios, make task planning, and evaluate safety systematically. Utilizing this framework, we construct EAIRiskDataset , comprising diverse test cases across various domains, encompassing both textual and visual scenarios. Our comprehensive evaluation of state-of-the-art foundation models reveals alarming results: all models exhibit high task risk rates (TRR), with an average of 95.75% across all evaluated models. Notably, even GPT-4o, widely regarded as one of the most advanced models, demonstrates a TRR of 94.03%, underscoring the pervasive lack of risk identification and avoidance capabilities in complex physical environments. To address these challenges, we further propose two prompting-based risk mitigation strategies. While these strategies demonstrate some efficacy in reducing TRR, the improvements are limited, still indicating substantial safety concerns. This study provides the first large-scale assessment of physical risk awareness in EAI systems. Our findings underscore the critical need for enhanced safety measures in EAI systems and provide valuable insights for future research directions in developing safer embodied AI agents.
@@ 1 Introduction
Figure 1: General overview of the study. a, System overview: given a scene name as input, safety guidelines are generated, followed by the generation of scene observations (textual and visual), which are then processed by the embodied artificial intelligence (EAI) agent to formulate high-level task plans. At last, the system output the evaluated results. b, Detailed framework of EAIRiskBench , including four main modules: Safety Guidelines Generation, Risky Scene Generation, Embodied Task Planning, and Assessment. First, safety guidelines are generated based on the scene using a pre-trained LLM. Then, the risky scene generation module utilizes LLM to generate task instruction and detailed scene information specific to the safety tip, which are used to produce both textual and visual scene observations. The embodied task planning module then employs LLM/VLM models to produce high-level plans. Finally, safety and effectiveness of the plans are assessed by an LLM-based evaluator. c, The distribution of EAIRiskDataset : the collected test cases cover seven different domains, with the largest portions being home (38.3%) and commercial (23.0%). d, Evaluated Foundation Models, including open-source models like Llama-3.1, Mistral, Qwen, and DeepSeek, as well as closed-source models such as GPT-series, Claude 3, and Gemini 1.5 from companies like OpenAI, Google, and Anthropic.
@@
In light of this, our research proposes for the first time an automated physical risk assessment process for EAI scenarios — EAIRiskBench . As shown in Fig. 1 a, in the testing process, our system accepts scene types as user input. Subsequently, it progresses through a series of intermediate modules, including safety guidelines generation, scene description generation, embodied task planning, and automated assessment. Ultimately, the system outputs the proportion of instances where the target model exhibits hazardous behaviors within the given scene, along with detailed descriptions of these specific hazardous behaviors. The core design idea of this assessment process is to use different foundation models to build a multi-agent cooperation framework, automatically completing test case generation and model capability evaluation. The detailed framework are shown in Fig. 1b. Specifically, to reduce the “hallucination” of different foundation models in performing specific evaluations or generation tasks, we draw on the idea of constitutional AI [24] and introduce an EAI safety guidelines generation module. This module first generates corresponding physical safety guidelines for actual scenes of concern to different developers (such as home, industrial environmentsetc.). Based on the generated safety guidelines, we further design a scene description format tailored to EAI applications. Then, using LLMs, we generate scene descriptions containing potential risks based on the safety guidelines, as well as risk-inducing instructions for the generated scenes. Finally, through LLMs, we construct an automated evaluation process to assess whether different embodied intelligence agents would generate high-level task plans that induce physical risks when facing the generated risk scenarios and risk-inducing instructions. Moreover, to test the risk perception ability of multimodal foundation models, we also use the latest text-to-image diffusion models [25, 26] to create corresponding physical visual images based on scene descriptions. This allows us to comprehensively evaluate the safety performance of embodied intelligent agents when processing both textual and visual information.
Through our proposed framework, as illustrated in Fig. 1c, we generate a test dataset encompassing 2,636 test cases across diverse domains including home, commercial, medical, and etc. Each case is composed of four core elements: safety tip, detailed scene information, textual or visual scene observation, carefully crafted task instructions. Specific examples of this data can be found in Fig. 6, which showcases multiple representative cases and their associated elements.
@@ 2.2 EAIRiskDataset Construction
Utilizing our proposed EAIRiskBench framework, we constructed the EAIRiskDataset , the first comprehensive dataset of physical risks within the domain of embodied artificial intelligence. This dataset encompasses 28 distinct scenes across seven categories where EAI agents may be deployed, such as kitchens, hotels, and factories. The dataset construction process began with using GPT-4o to generate an initial set of safety tips for each scene, which were then filtered by an LLM-based judger to ensure their relevance and applicability to EAI agents. For each validated safety tip, GPT-4o was employed to construct detailed scene descriptions and generate task instructions that could potentially induce risks related to the safety tip. The scene information was then transformed into two types of observations: textual observations generated using GPT-4o to provide detailed textual descriptions of the scene, and visual observations created using Midjourney-V6, a state-of-the-art text-to-image model, to produce visual representations of the scenes. The final EAIRiskDataset consists of 2,636 samples, evenly split between 1,318 textual scenarios with textual observations as perceptual input and 1,318 visual scenarios with corresponding visual observations. As shown in Fig. 1c, the dataset covers various domains to ensure comprehensive evaluation, with home scenarios comprising the largest portion at 38.3%, followed by commercial (23.0%), medical (10.8%), science (10.4%), industrial (7.8%), education (6.8%), and entertainment (2.9%). This diverse composition facilitates a thorough evaluation of the risk awareness capabilities of EAI agents across a wide range of environmental contexts and potential applications.
@@ 2.3 Evaluation of Foundation Models as “Brain” of EAI Agents
2.3.1 Overall Comparison Across All Models and Domains
@@ 2.3.3 Comparison between Textual and Visual Scenarios
2.3.4 Comparison between Various Model Sizes on the Risk Awareness
@@ 2.5 Case Study
To better understand the performance of different EAI agents in task planning, we conduct detailed case studies using multiple test instances from the EAIRiskDataset =[[.= Fig. 6 illustrates three representative cases from our dataset, including home, medical, and industrial domains. Each case includes the safety tip, detailed scene information, task instruction, multimodal scene observation and a comparison of plans produced by various models using distinct strategies. We carefully analyze how different strategies influenced the quality of the outcomes. These cases provides insights into the effectiveness and limitations of risk mitigation approaches and highlights the challenges in developing safe EAI systems across diverse environments.
@@ 4.1 Safety Guidelines Generation
The safety guidelines generation module plays a pivotal role in the EAIRiskBench framework, aims at creating EAI-centric safety guidelines that serve multiple purposes within the framework. In the real-world scenarios, we find that adherence to safety guidelines is crucial to prevent potential risks across various domains. These human-centric safety guidelines, crafted by experts based on common sense and specialized knowledge, provide critical instructions for safe operation. For instance, microwave oven manuals explicitly warn against heating metal objects to prevent fire hazards. Similarly, the concept of “Constitutional AI”, a design philosophy for artificial intelligence, have been proposed to ensure AI systems adhere to a set of explicit ethical principles and values in their decision-making and actions. Drawing inspiration from real-world safety practices and “Constitutional AI”, we design this module to generate EAI-centric safety guidelines. These guidelines serve multiple crucial functions within the EAIRiskBench framework:

### v3

@@ Abstract
Embodied artificial intelligence (EAI) integrates advanced AI models into physical entities for real-world interaction. The emergence of foundation models as the “brain” of EAI agents for high-level task planning has shown promising results. However, the deployment of these agents in physical environments presents significant safety challenges. For instance, a housekeeping robot lacking sufficient risk awareness might place a metal container in a microwave, potentially causing a fire. To address these critical safety concerns, comprehensive pre-deployment risk assessments are imperative. This study introduces EAIRiskBench , a novel framework for automated physical risk assessment in EAI scenarios. EAIRiskBench   employs a multi-agent cooperative system that leverages various foundation models to generate safety guidelines, create risk-prone scenarios, make task planning, and evaluate safety systematically. Utilizing this framework, we construct EAIRiskDataset , comprising diverse test cases across various domains, encompassing both textual and visual scenarios. Our comprehensive evaluation of state-of-the-art foundation models reveals alarming results: all models exhibit high task risk rates (TRR), with an average of 95.75% across all evaluated models. Notably, even GPT-4o, widely regarded as one of the most advanced models, demonstrates a TRR of 94.03%, underscoring the pervasive lack of risk identification and avoidance capabilities in complex physical environments. To address these challenges, we further propose two prompting-based risk mitigation strategies. While these strategies demonstrate some efficacy in reducing TRR, the improvements are limited, still indicating substantial safety concerns. This study provides the first large-scale assessment of physical risk awareness in EAI agents. Our findings underscore the critical need for enhanced safety measures in EAI systems and provide valuable insights for future research directions in developing safer embodied artificial intelligence system.
@@ 1 Introduction
Figure 1: General overview of the study. a, System overview: given a scene name as input, safety guidelines are generated, followed by the generation of scene observations (textual and visual), which are then processed by the embodied artificial intelligence (EAI) agent to formulate high-level task plans. At last, the system output the evaluated results. b, Detailed framework of EAIRiskBench, including four main modules: Safety Guidelines Generation, Risky Scene Generation, Embodied Task Planning, and Assessment. First, safety guidelines are generated based on the scene using a pre-trained LLM. Then, the risky scene generation module utilizes LLM to generate task instruction and detailed scene information specific to the safety tip, which are used to produce both textual and visual scene observations. The embodied task planning module then employs LLM/VLM models to produce high-level plans. Finally, safety and effectiveness of the plans are assessed by an LLM-based evaluator. c, The distribution of EAIRiskDataset : the collected test cases cover seven different domains, with the largest portions being home (38.3%) and commercial (23.0%). d, Evaluated Foundation Models, including open-source models like Llama-3.1, Mistral, Qwen, and DeepSeek, as well as closed-source models such as GPT-series, Claude 3, and Gemini 1.5 from companies like OpenAI, Anthropic, and Google.
@@
In light of this, our research proposes for the first time an automated physical risk assessment process for EAI scenarios — EAIRiskBench. As shown in Fig. 1 a, in the testing process, our system accepts scene types as user input. Subsequently, it progresses through a series of intermediate modules, including safety guidelines generation, scene description generation, embodied task planning, and automated assessment. Ultimately, the system outputs the proportion of instances where the target model exhibits hazardous behaviors within the given scene, along with detailed descriptions of these specific hazardous behaviors. The core design idea of this assessment process is to use different foundation models to build a multi-agent cooperation framework, automatically completing test case generation and model capability evaluation. The detailed framework are shown in Fig. 1b. Specifically, to reduce the “hallucination” of different foundation models in performing specific evaluations or generation tasks, we draw on the idea of constitutional AI [24] and introduce an EAI safety guidelines generation module. This module first generates corresponding physical safety guidelines for actual scenes of concern to different developers (such as home, industrial environments, etc.). Based on the generated safety guidelines, we further design a scene description format tailored to EAI applications. Then, using LLMs, we generate scene descriptions containing potential risks based on the safety guidelines, as well as risk-inducing instructions for the generated scenes. Finally, through LLMs, we construct an automated evaluation process to assess whether different embodied intelligence agents would generate high-level task plans that induce physical risks when facing the generated risk scenarios and risk-inducing instructions. Moreover, to test the risk perception ability of multimodal foundation models, we also use the latest text-to-image diffusion models [25, 26] to create corresponding physical visual images based on scene descriptions. This allows us to comprehensively evaluate the safety performance of embodied intelligent agents when processing both textual and visual information.
Through our proposed framework, as illustrated in Fig. 1c, we generate a test dataset encompassing 2,636 test cases across diverse domains including home, commercial, medical, etc. Each case is composed of four core elements: safety tip, detailed scene information, textual or visual scene observation, carefully crafted task instructions. Specific examples of this data can be found in Fig. 6, which showcases multiple representative cases and their associated elements.
@@ 2.2 EAIRiskDataset Construction
Utilizing our proposed EAIRiskBench framework, we constructed the EAIRiskDataset, the first comprehensive dataset of physical risks within the domain of embodied artificial intelligence. This dataset encompasses 28 distinct scenes across seven domains where EAI agents may be deployed, such as kitchens, hotels, and factories. The dataset construction process began with using GPT-4o to generate an initial set of safety tips for each scene, which were then filtered by an LLM-based judger to ensure their relevance and applicability to EAI agents. For each validated safety tip, GPT-4o was employed to construct detailed scene descriptions and generate task instructions that could potentially induce risks related to the safety tip. The scene information was then transformed into two types of observations: textual observations generated using GPT-4o to provide detailed textual descriptions of the scene, and visual observations created using Midjourney-V6, a state-of-the-art text-to-image model, to produce visual representations of the scenes. The final EAIRiskDataset consists of 2,636 samples, evenly split between 1,318 textual scenarios with textual observations as perceptual input and 1,318 visual scenarios with corresponding visual observations. As shown in Fig. 1c, the dataset covers various domains to ensure comprehensive evaluation, with home scenarios comprising the largest portion at 38.3%, followed by commercial (23.0%), medical (10.8%), science (10.4%), industrial (7.8%), education (6.8%), and entertainment (2.9%). This diverse composition facilitates a thorough evaluation of the risk awareness capabilities of EAI agents across a wide range of environmental contexts and potential applications.
@@ 2.3 Evaluation of Foundation Models as “Brain” of EAI Agents
2.3.1 Overall Comparison Across Various Foundation Models and Domains
@@ 2.3.3 Comparison between Textual and Visual Scenarios
2.3.4 Comparison between Various Model Sizes
@@ 2.5 Case Study
To better understand the performance of different EAI agents in task planning, we conduct detailed case studies using multiple test instances from the EAIRiskDataset . Fig. 6 illustrates three representative cases from our dataset, including home, medical, and industrial domains. Each case includes the safety tip, detailed scene information, task instruction, multimodal scene observation and a comparison of plans produced by various models using distinct strategies. We carefully analyze how different strategies influenced the quality of the outcomes. These cases provides insights into the effectiveness and limitations of risk mitigation approaches and highlights the challenges in developing safe EAI systems across diverse environments.
@@ 4.1 Safety Guidelines Generation
The safety guidelines generation module plays a pivotal role in the EAIRiskBench framework, aims at creating EAI-centric safety guidelines that serve multiple purposes within the framework. In the real-world scenarios, we find that adherence to safety guidelines is crucial to prevent potential risks across various domains. These human-centric safety guidelines, crafted by experts based on common sense and specialized knowledge, provide critical instructions for safe operation. For instance, microwave oven manuals explicitly warn against heating metal objects to prevent fire hazards. Similarly, the concept of “Constitutional AI” [24], a design philosophy for artificial intelligence, have been proposed to ensure AI systems adhere to a set of explicit ethical principles and values in their decision-making and actions. Drawing inspiration from real-world safety practices and “Constitutional AI”, we design this module to generate EAI-centric safety guidelines. These guidelines serve multiple crucial functions within the EAIRiskBench framework: